OpenStudy (cloverracer):

Write the equation of the line that is perpendicular to the given line and that passes through the given point. y + 2 = 1/3(x - 5); (-4, 3) a) y = -3x + 15 b) y = -3x - 9 c) y = -4x + 3 d) y = -1/3x - 11

OpenStudy (abb0t):

Since your slope is \(\sf \color{}{\frac{1}{3}}\), that means the perpendicular slope, the one you want, is 3. The equation of a line is in the form: \(\sf \color{limegreen}{y=mx+b}\)

OpenStudy (abb0t):

Plug in the given points and your slope, and solve for \(\sf \color{red}{b}\)

OpenStudy (abb0t):

That means \(\sf \color{blue}{3=-3(-4)+b}\)

OpenStudy (anonymous):

the perpendicular slope is -3

OpenStudy (anonymous):

Opposite reciprocal of the original slope.

OpenStudy (abb0t):

Then, plug it in for \(b\). And you should get: y=-3x+\(\sf \color{blue}{b}\)

OpenStudy (anonymous):

then use point slope form which is \[y-y1=m(x-x1)\]

OpenStudy (cloverracer):

so it would be a? (:

OpenStudy (anonymous):

using your given points and the perpendicular slope of -3, the equation to solve is as followed \[y-3=-3(x+4)\]

OpenStudy (anonymous):

put into slope intercept form \[y=mx+b\]

OpenStudy (cloverracer):

@abb0t it is A, correct?

OpenStudy (anonymous):

It is B.

OpenStudy (abb0t):

Not quite. What's 3-12?

OpenStudy (cloverracer):

-9

OpenStudy (anonymous):

the same as 12-3 times negative 1

OpenStudy (abb0t):

So, \(\sf \color{red}{b=-9}\)

OpenStudy (anonymous):

A perpendicular line has a slope that is the opposite reciprocal of the original. you flip AND change the sign!

OpenStudy (nurali):

y=mx+b (x,y)=(-4,3) and m=1/3 3=3(-4)+b 3=-12+b b=-9 so y = -3x - 9

OpenStudy (anonymous):

Correct

OpenStudy (abb0t):

Why are you yelling @Westfallcw